A Sumon of tk. 53 is divided among A, B, C in such a way that A gets tk. 7 more than that of B gets & B gets tk 8 more than C gets. The ratio of their shares is what?
Answer by stanbon(75887)

A Sumon of tk. 53 is divided among A, B, C in such a way that A gets tk. 7 more than that of B gets & B gets tk 8 more than C gets. The ratio of their shares is what?
===============
a + b + c = 53
a - b = 7
c - b = -8
----------------------
Substitute for "a" and for "c":: solve for "b"::
b+7 + b + b-8 = 53
-----
3b = 54
b = 18 shares
-----
a = b+7 = 25 shares
-----
c = b-8 = 10 shares
------
Cheers,
Stan H.
-------------